OpenStudy (anonymous):

who did the functions unit test 8th grade pre algebra that goes to connections??? WILL MEDAL AND FAN

OpenStudy (anonymous):

I already did pre-algebra...What is the question?

OpenStudy (anonymous):

weell actually its a whole test

OpenStudy (anonymous):

I will help you find the answers because I will not help you cheat...I will tell you how to find it and you can use a calculator if needed

OpenStudy (anonymous):

i might give you some answers but few

OpenStudy (anonymous):

I WILL HELP YOU.....what are the questions

OpenStudy (anonymous):

which functioīi rule represents the data in the table below. input (x)=1-2-3-4-5 output (y)=9-15-21-27-33

OpenStudy (anonymous):

ill cheat 4 u

OpenStudy (anonymous):

can anyone give me the answers to these questions please

OpenStudy (anonymous):



OpenStudy (anonymous):

for #6

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

I meant the last one

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

really because i guessed

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

b

OpenStudy (anonymous):

B

OpenStudy (anonymous):

right

OpenStudy (anonymous):

c for the 1st one

OpenStudy (anonymous):

and also c for the second one

OpenStudy (anonymous):

that one is tough..what do you think?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

a I think because it is a negative slope

OpenStudy (anonymous):

This one I do not know...Its up to you this time

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):

Just try and graph it and solve

OpenStudy (anonymous):

yea i donnt know

OpenStudy (anonymous):

Ill try again...sorry

OpenStudy (anonymous):

It doesnt look like C...